# DrugAgent: Explainable Drug Repurposing Agent with Large Language Model-based Reasoning

### v1

@@ 4.3 Implementation Details
4.4 Case Study
Table 1: Case study 1: known drug-target interaction
User Input
Problem: I have designed a drug repurposing study and hope you can help me predict whether this drug works for the target. Features contain (1) drug: Topotecan; (2) target: TOP1;
Coordinate (Coordinator Agent): Coordinate the problem
Thought: Decompose a complex problem into several easier subproblems.
Subproblem 1 (AI Agent): Predict the DTI score using pre-trained DeepPurpose model. Subproblem 2 (Search Agent): Analyze DTI data using search methods and scientific literature. Subproblem 3 (Knowledge Graph Agent): Assess the relationship between Topotecan and TOP1 using knowledge graphs.
Solving Subproblems:
Subproblem 1 (AI Agent): The ML-based DTI score for Topotecan and TOP1 is calculated as 7.649889945983887. This high score suggests a strong potential interaction. Subproblem 2 (Search Agent): Analysis of scientific literature yields a search-based DTI score of 0.27. This low score might indicate limited direct evidence in published research. Subproblem 3 (Knowledge Graph Agent): The knowledge graph analysis results in a score of 1, indicating a known connection between Topotecan and TOP1 in existing databases.
Reasoning (Reasoning Agent): Integrating multi-source information
1. ML Prediction: The high ML-based score (7.65) suggests that the structural and chemical properties of Topotecan are highly compatible with TOP1 binding, indicating a strong potential for interaction. 2. Literature Evidence: The low search-based score (0.27) implies that while there might be some mentions of Topotecan and TOP1 in literature, the direct evidence of their interaction is not extensively documented or studied. 3. Known Connections: The knowledge graph score of 1 confirms that there is an established connection between Topotecan and TOP1 in curated databases. 4. Background Information: Topotecan is a known topoisomerase I inhibitor, and TOP1 (DNA topoisomerase I) is its primary target. It is used in cancer treatment, particularly for ovarian and small-cell lung cancers.
Conclusion: Merging the scores with $\alpha=0.3$ and $\beta=0.3$ yields a final drug repurposing score of 4.059966689596993. This moderate-to-high score, combined with the known mechanism of action, strongly suggests that Topotecan is effective against TOP1. (The ground truth confirms Topotecan’s effectiveness against TOP1)
Table 2: Case study 2: unknown, but might have some relationship drug-target interaction
User Input
Problem: I have designed a drug repurposing study and hope you can help me predict whether this drug works for the target. Features contain (1) drug: Topotecan; (2) target: SLFN11;
Coordinate (Coordinator Agent): Coordinate the problem  Thought: Decompose a complex problem into several easier subproblems.   Subproblem 1 (AI Agent): Predict the DTI score using pre-trained DeepPurpose model. Subproblem 2 (Search Agent): Analyze DTI data using search methods and scientific literature. Subproblem 3 (Knowledge Graph Agent): Assess the relationship between Topotecan and SLFN11 using knowledge graphs.
Solving Subproblems:
Subproblem 1 (AI Agent): The ML-based DTI score for Topotecan and SLFN11 is calculated as 7.363409519195557. This high score suggests a strong potential interaction. Subproblem 2 (Search Agent): Analysis of scientific literature yields a search-based DTI score of 0.33. This low-to-moderate score might indicate some evidence of interaction in published research, but not extensively documented. Subproblem 3 (Knowledge Graph Agent): The knowledge graph analysis results in a score of 0.7213475204444817, indicating a moderate connection between Topotecan and SLFN11 in existing databases.
Reasoning (Reasoning Agent): Integrating multi-source information
1. ML Prediction: The high ML-based score (7.36) suggests that the structural and chemical properties of Topotecan are potentially compatible with SLFN11 interaction, indicating a strong potential for some form of relationship. 2. Literature Evidence: The low-to-moderate search-based score (0.33) implies that there is some mention of Topotecan and SLFN11 in literature, suggesting a potential relationship, though not as extensively studied as more established DTIs. 3. Known Connections: The knowledge graph score of 0.72 indicates a moderate level of established connection between Topotecan and SLFN11 in curated databases, suggesting some recognized relationships in the scientific community. 4. Background Information: While Topotecan is primarily known as a topoisomerase I inhibitor, its potential relationship with SLFN11 could be indirect, possibly through DNA damage response pathways that SLFN11 is involved in.  Conclusion: Merging the scores with $\alpha=0.3$ and $\beta=0.3$ yields a final drug repurposing score of 3.1383244829891226. This moderate-to-high score suggests that there might be a meaningful relationship between Topotecan and SLFN11, possibly mediated indirectly through DNA damage response pathways. Further investigation would be warranted to clarify the nature and extent of this potential interaction.
Table 3: Case study 3: unknown drug-target interaction
User Input
Problem: I have designed a drug repurposing study and hope you can help me predict whether this drug works for the target. Features contain (1) drug: Topotecan; (2) target: SLC26A4;
Coordinate (Coordinator Agent): Coordinate the problem  Thought: Decompose a complex problem into several easier subproblems.   Subproblem 1 (AI Agent): Predict the DTI score using pre-trained DeepPurpose model. Subproblem 2 (Search Agent): Analyze DTI data using search methods and scientific literature. Subproblem 3 (Knowledge Graph Agent): Assess the relationship between Topotecan and SLC26A4 using knowledge graphs.
Solving Subproblems:
Subproblem 1 (AI Agent): The ML-based DTI score for Topotecan and SLC26A4 is calculated as 7.609444618225098. This high score suggests a strong potential interaction, based on structural and chemical properties. Subproblem 2 (Search Agent): Analysis of scientific literature yields a search-based DTI score of 0.07. This very low score indicates extremely limited or no direct evidence of interaction in published research. Subproblem 3 (Knowledge Graph Agent): The knowledge graph analysis results in a score of 0.7213475204444817, indicating a moderate connection between Topotecan and SLC26A4 in existing databases.
Reasoning (Reasoning Agent): Integrating multi-source information
1. ML Prediction: The high ML-based score (7.61) suggests that the structural and chemical properties of Topotecan might be compatible with SLC26A4 binding. However, this prediction is based solely on molecular features and does not consider biological context. 2. Literature Evidence: The extremely low search-based score (0.07) strongly implies that there is little to no documented evidence of direct interaction between Topotecan and SLC26A4 in scientific literature. 3. Known Connections: The moderate knowledge graph score (0.72) suggests some indirect connections between Topotecan and SLC26A4 in curated databases, possibly through shared pathways or related molecular processes. 4. Background Information: Topotecan is primarily known as a topoisomerase I inhibitor used in cancer treatment. SLC26A4, on the other hand, is a protein involved in iodide transport in the inner ear and thyroid. These functions appear unrelated, which aligns with the low literature evidence score.
Conclusion: Merging the scores with $\alpha=0.3$ and $\beta=0.3$ yields a final drug repurposing score of 2.3765816855913298. This moderate score, primarily driven by the ML prediction and knowledge graph connection, suggests a potential for interaction between Topotecan and SLC26A4. However, the very low literature evidence score indicates that this interaction, if it exists, is not well-studied or documented. The discrepancy between the ML prediction and literature evidence highlights the need for careful interpretation and further experimental validation. Given the known functions of Topotecan and SLC26A4, any potential interaction would likely be indirect or through previously unrecognized mechanisms. Further investigation would be necessary to understand the nature and relevance of any possible interaction between Topotecan and SLC26A4.
We analyze three realistic case studies focusing on drug repurposing for Topotecan, targeting different proteins: TOP1, SLFN11, and SLC26A4 (Tables 1, 2, and 3). Topotecan inhibits TOP1, preventing DNA religation and causing cell death in rapidly dividing cancer cells (Schmidt et al., 2001). While SLFN11 is not the direct target of Topotecan, it has been associated with Topotecan efficacy in some studies (Raynaud et al., 2023). In contrast, SLC26A4, a gene encoding pendrin (an anion exchanger protein), is not directly related to Topotecan’s mechanism of action or efficacy. These cases represent a spectrum of DTIs, from well-established to potentially novel and unlikely connections. For each case, the user initiates the process by providing the drug name (Topotecan) and the target protein. The Coordinator Agent then decomposes the complex problem into three subproblems, handled by specialized agents: the AI Agent, the Search Agent, and the Knowledge Graph Agent. Case 1 (Topotecan-TOP1) represents a known strong interaction. The AI Agent calculates a high score of 7.65, indicating strong potential interaction based on structural properties. The Search Agent yields a relatively low score of 0.27, possibly due to the well-established nature of this interaction not requiring extensive new studies. The Knowledge Graph Agent confirms this with a score of 1, indicating a known connection in existing databases. Case 2 (Topotecan-SLFN11) explores a less understood but potentially relevant interaction. The AI Agent predicts a high score of 7.36, suggesting structural compatibility. The Search Agent’s score of 0.33 indicates some literature evidence, while the Knowledge Graph Agent’s score of 0.72 suggests moderate established connections. Case 3 (Topotecan-SLC26A4) investigates an unlikely interaction. Despite a high ML score of 7.61, the Search Agent’s very low score of 0.07 indicates minimal literature evidence. The Knowledge Graph Agent’s score of 0.72 suggests some indirect connections, highlighting the system’s ability to detect potential novel interactions while also recognizing the need for careful interpretation. In each case, the Reasoning Agent integrates this multi-source information, considering the ML-based structural compatibility, literature-based evidence, known connections, and relevant background information. The final drug repurposing scores are calculated using weights $\alpha=0.3$ and $\beta=0.3$, resulting in scores of 4.06, 3.14, and 2.38 for cases 1, 2, and 3, respectively. These case studies demonstrate the multi-agent system’s ability to:
  1. 1.
    Handle diverse scenarios, from known interactions to potentially novel discoveries.
  2. 2.
    Leverage and integrate various data sources and analytical methods.
  3. 3.
    Provide interpretable results with detailed reasoning processes.
  4. 4.
    Recognize system limitations, especially when predictions conflict with existing evidence.
  5. 5.
    Offer practical insights for drug repurposing research across different levels of prior knowledge.
  6. 6.
    Highlight areas requiring further investigation or experimental validation.
This comprehensive approach not only confirms known interactions but also suggests potential new avenues for research, showcasing the system’s value in accelerating and refining the drug repurposing process.
5 Discussion
Our study presents a novel multi-agent system for drug repurposing that integrates machine learning, knowledge graphs, and literature search. This approach offers more robust predictions by leveraging diverse data sources and analytical methods.
The system’s strength lies in its collaborative approach, which combines each agent’s specialized capabilities to evaluate complex DTIs comprehensively. The weighted integration method allows for flexible adjustment of different prediction methods, enhancing overall accuracy.
However, limitations exist. The system still relies on human expertise for initial setup, limiting its scalability. It also lacks autonomous knowledge updating capabilities to keep pace with rapidly evolving pharmacological research. Furthermore, the current model does not adequately account for individual patient characteristics or drug combination effects.
Future research should focus on:
  • Integrating autonomous knowledge updating mechanisms to keep the system current with the latest pharmacological research;
  • Enhancing the system’s ability to predict drug efficacy and handle complex drug combinations;
  • Expanding the system’s applicability to a broader range of pharmaceutical tasks by incorporating existing models:
    • Adapting the framework for Drug Response Prediction using models like drGAT (Inoue et al., 2024b);
    • Extending to DTI prediction tasks, leveraging the system’s existing knowledge graph and AI capabilities;
    • Developing a more flexible prompting system to allow easy adaptation to various pharmaceutical tasks without major architectural changes, such as drug synergy prediction (Huang et al., 2022), drug property prediction (Xu et al., 2024), adverse drug reaction prediction (Chen et al., 2024a) or drug design (Fu et al., 2021a; 2022a);
  • Automating the preprocessing of complex data types, such as scRNA-seq data:
    • Implementing multiple preprocessing functions including Imputation (Inoue et al., 2024a; Inoue, 2024), Quality Control (McCarthy et al., 2017; Lu et al., 2023), and batch effect correction (Li et al., 2020; Fu et al., 2024; Haghverdi et al., 2018);
    • Developing an optimization framework to automatically select and apply the most appropriate preprocessing methods for given datasets;
    • Integrating these preprocessing capabilities seamlessly into the existing multi-agent system;
  • Validating the expanded system’s predictions through experimental studies and clinical trials to ensure real-world applicability across multiple tasks and data types (Lu et al., 2024a; Fu et al., 2023).
Experimental Setup
@@ 5 Discussion
Aknoeledge
We thank the DGIdb and DrugBank team for granting permission to use their logo in this publication.
Appendix A Appendix
A.1 Weight Optimization for Score Integration
To determine the optimal weights for our score integration process, we employed a simple optimization method. We utilized a dataset comprising 3,347 drug-target pairs, each containing scores from our three specialized agents (AI, Knowledge Graph, and Search) along with corresponding ground truth interaction scores from the BindingDB dataset. We formulated the weight optimization as a constrained least squares problem:
$\underset{\mathbf{x}}{\arg\min}$ $||\mathbf{Ax}-\mathbf{b}||^{2}_{2}$ subject to $\sum{i=1}^{3}x_{i}=1$ $x_{i}\geq 0,\quad i=1,2,3$ (3)
where $\mathbf{A}$ is the matrix of agent scores, $\mathbf{x}$ is the weight vector we are optimizing, and $\mathbf{b}$ is the vector of ground truth scores. The optimization was performed using non-negative least squares (NNLS) followed by Sequential Least Squares Programming (SLSQP) to ensure the weights sum to 1. This approach allows us to determine the relative importance of each agent’s score in predicting the true DTI strength. The resulting weights were then used in our merged DTI score calculation:
$S_{\text{merged}}=x_{1}S_{\text{AI}}+x_{2}S_{\text{KG}}+x_{3}S_{\text{Search}},$ (4)
where $x_{1}$, $x_{2}$, and $x_{3}$ represent the optimized weights for the AI Agent, Knowledge Graph Agent, and Search Agent, respectively. This data-driven weight optimization approach ensures that our final predictions leverage the strengths of each agent while accounting for their relative performance on a large set of known DTIs.
A.2 Detailed External Tool Definitions
@@
A.2.1 AI Agent Implementation
The AI Agent utilizes machine learning models to predict DTIs. Its core function, get_ml_dti_score, takes a drug name and a target name as input and returns a float value representing the predicted interaction score.
1# AI Agent
2def get_ml_dti_score(name: str, target_name: str) -> float:
3 target_sequence = get_target_sequence(target_name)
4 net = models.model_pretrained(model="MPNN_CNN_BindingDB")
5
6 X_repurpose, drug_name, drug_cid = load_broad_repurposing_hub(
7 SAVE_PATH
8 )
9 idx = drug_name == name
10 if not any(idx):
11 print(f"Logging: Drug ’{name}’ not found.")
12 return None
13
14 res = models.virtual_screening(
15 X_repurpose[idx], [target_sequence], net,
16 drug_name[idx], [target_name]
17 )
18 return res[0]
@@ 15 X_repurpose[idx], [target_sequence], net,
A.2.2 Search Agent Implementation
@@
1# Search Agent
2def google_search(query: str, num_results: int = 10) -> List[Dict[str, str]]:
3 # ... [implementation details]
4
5def _parse_search_results(soup: BeautifulSoup) -> List[Dict[str, str]]:
6 # ... [implementation details]
7
8def calculate_dti_score(search_results: List[Dict[str, str]],
9 drug_name: str, target_name: str) -> float:
10 # ... [implementation details]
11
12def _calculate_individual_score(result: Dict[str, str], drug_name: str,
13 target_name: str, positive_keywords: List[str],
14 strong_keywords: List[str]) -> int:
15 # ... [implementation details]
16
17def analyze_dti(name: str, target_name: str) -> float:
18 search_results = google_search(f"{name} {target_name} interaction")
19 dti_score = calculate_dti_score(search_results, name, target_name)
20 return dti_score
@@ A.2.3 KG Agent Implementation
1# KG Agent
2def calculate_dti_score(kg, drug, target):
3 if drug not in kg.graph or target not in kg.graph:
4 return 0 # Return 0 if the drug or target is not in the knowledge graph
5
6 hops = kg.shortest_path(drug, target)
7 if hops == -1:
8 return 0 # No relationship
9 elif hops == 1:
10 return 1 # Direct connection
11 else:
12 return 1 / (np.log1p(hops)) # Logarithm-based score
13
14def load_kg(file_path):
15 with open(file_path, "rb") as f:
16 kg = pickle.load(f)
17 return kg
18
19def get_kg_dti_score(name: str, target_name: str) -> float:
20 kg = load_kg("../data/knowledge_graph.pkl")
21 score = calculate_dti_score(kg, name, target_name)
22 return score

### v2

@@ A.2 Implementation Details
Software and Hardware Configuration
@@
A.3 Detailed External Tool Definitions
@@
A.3.1 AI Agent Implementation
The AI Agent utilizes machine learning models to predict DTIs. Its core function, get_AI_score, takes a drug name and a target name as input and returns a float value representing the predicted interaction score.
# AI Agent
def get_ml_dti_score(name: str, target_name: str) -> float:
target_sequence = get_target_sequence(target_name)
net = models.model_pretrained(model="MPNN_CNN_BindingDB")
X_repurpose, drug_name, drug_cid = load_broad_repurposing_hub(
SAVE_PATH
idx = drug_name == name
if not any(idx):
print(f"Logging: Drug ’{name}’ not found.")
return None
res = models.virtual_screening(
X_repurpose[idx], [target_sequence], net,
drug_name[idx], [target_name]
return res[0]
@@ A.3.1 AI Agent Implementation
A.3.2 Search Agent Implementation
@@
# Search Agent
def google_search(query: str, num_results: int = 10) -> List[Dict[str, str]]:
# ... [implementation details]
def _parse_search_results(soup: BeautifulSoup) -> List[Dict[str, str]]:
# ... [implementation details]
def calculate_dti_score(search_results: List[Dict[str, str]],
drug_name: str, target_name: str) -> float:
# ... [implementation details]
def _calculate_individual_score(result: Dict[str, str], drug_name: str,
target_name: str, positive_keywords: List[str],
strong_keywords: List[str]) -> int:
# ... [implementation details]
def analyze_dti(name: str, target_name: str) -> float:
search_results = google_search(f"{name} {target_name} interaction")
dti_score = calculate_dti_score(search_results, name, target_name)
return dti_score
@@ A.3.3 KG Agent Implementation
# KG Agent
def calculate_dti_score(kg, drug, target):
if drug not in kg.graph or target not in kg.graph:
return 0 # Return 0 if the drug or target is not in the knowledge graph
hops = kg.shortest_path(drug, target)
if hops == -1:
return 0 # No relationship
elif hops == 1:
return 1 # Direct connection
else:
return 1 / (np.log1p(hops)) # Logarithm-based score
def load_kg(file_path):
with open(file_path, "rb") as f:
kg = pickle.load(f)
return kg
def get_kg_dti_score(name: str, target_name: str) -> float:
kg = load_kg("../data/knowledge_graph.pkl")
score = calculate_dti_score(kg, name, target_name)
return score
@@
Appendix B Experimental Setup
B.1 Dataset and Evaluation Metrics
Our regression task involved predicting PKD (negative logarithm of the dissociation constant Kd) values from the BindingDB database, given the target protein’s amino acid sequence and the drug compound’s SMILES string (Simplified Molecular-Input Line-Entry System, a line notation for encoding molecular structures) (Weininger, 1988). This task is crucial for understanding Drug-Target Interactions (DTIs) and has significant implications for drug discovery and development processes (Huang et al., 2022).
To comprehensively evaluate our model’s performance, we employed a suite of seven complementary metrics: Mean Squared Error (MSE), Mean Absolute Percentage Error (MAPE), Mean Absolute Error (MAE), R-squared (R2) Score, Explained Variance, Maximum Error, and Correlation. These metrics collectively assess various aspects of our predictions: MSE and MAE provide measures of the average prediction error, with MSE being more sensitive to large errors due to its quadratic nature. MAPE offers insight into the relative size of prediction errors. The R2 score and Explained Variance evaluate the model’s capacity to capture the underlying variance in the data. Maximum Error highlights the worst-case prediction scenario, crucial for understanding the model’s limitations. Lastly, Correlation assesses the strength and direction of the relationship between predicted and actual PKD values.
This comprehensive set of metrics allows us to thoroughly assess the accuracy of our predictions, the model’s capacity to capture underlying patterns in DTIs, and its consistency across different scenarios. Such a multi-faceted evaluation is essential for validating the model’s performance and identifying areas for potential improvement in the context of PKD score prediction for drug-target interactions. Due to space limitation, implementation details are provided in Appendix (Section A.2).
B.2 Procedure